OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±(5/6) x.

OpenStudy (anonymous):

@phi @robz8 ?!?!

OpenStudy (anonymous):

we can do this i am sure first off the center is \((0,0)\) so you know it looks like \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

second, since the vertices are at \((0,\pm10)\) we know \(a=10\) so we get to \[\frac{x^2}{10^2}-\frac{y^2}{b^2}=1\] so all we need now it \(b\)

OpenStudy (anonymous):

a little weird at this point though you sure it is \(y=\frac{5}{6}x\) right?

OpenStudy (anonymous):

\[\frac{b}{a}=\frac{5}{6}\] so \[\frac{b}{10}=\frac{5}{6}\] giving \(b=\frac{50}{6}\) hmm

OpenStudy (anonymous):

ok if that is what it is, then that is what it is \[\frac{x^2}{100}-\frac{36y^2}{2500}=1\] lets check it

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=hyperbola+ \frac{x^2}{100}-\frac{36y^2}{2500}%3D1 yup looks good to me

OpenStudy (anonymous):

sorry about that, thats what i thought. THANK YOU

OpenStudy (anonymous):

@satellite73